Use synthetic division to divide f(x) by X-c then write f(x) in the form f(x)=(x - c)9(x) +r. f(x)=x* + x° +5x; x+2
Answer #1

Solia f(u) = x + x3 +5k Devide by (x+2) x = -2 0 5 o 1 -2 1 O -2 2 -4 -2 2 1 - 1 11-2 PCX)= x3_r?+2x+1 { v=-2 f(x) = (x+2)(